Kenosha Winter Services is a small, family-owned snow-removal business. For its services, the company has always charged a flat fee per hundred square metres of snow removal. The current fee is $11.55 per hundred square metres. However, there is some question about whether the company is actually making any money on jobs for some customers—particularly those located on more remote properties that require considerable travel time. The owner’s daughter, home from school for the summer, has suggested investigating this question using ABC. After some discussion, a simple system consisting of four activity cost pools seemed to be adequate. The activity cost pools and their activity measures appear below:

  Activity Cost Pool Activity Measure Activity for the Year
  Snow removal Square metres cleaned (00s) 40,055 hundred square metres
  Travel to jobs Kilometres driven 20,625 kilometres
  Job support Number of jobs 250 jobs
  Other (costs of idle capacity and
     organization-sustaining costs)
None NA

The total cost of operating the company for the year is $442,500, which includes the following costs:

  
  Wages $ 171,000
  Supplies 43,500
  Snow removal equipment depreciation 23,500
  Vehicle expenses 43,500
  Office expenses 67,000
  President’s compensation 94,000
  
  Total cost $ 442,500
  

Resource consumption is distributed across the activities as follows:

Distribution of Resource Consumption across
Activity Cost Pools
Snow
Removal
Travel
to Jobs
Job
Support
Other Total
  Wages 80 % 10 % 0 % 10 % 100 %
  Supplies 100 % 0 % 0 % 0 % 100 %
  Snow removal equipment depreciation 85 % 0 % 0 % 15 % 100 %
  Vehicle expenses 0 % 65 % 0 % 35 % 100 %
  Office expenses 0 % 0 % 45 % 55 % 100 %
  President’s compensation 0 % 0 % 35 % 65 % 100 %

Job support consists of receiving calls from potential customers at the home office, scheduling jobs, billing, resolving issues, and so on.

Required:

1. Prepare the first-stage allocation of costs to the activity cost pools. (Do not leave any empty spaces; input a 0 wherever it is required.)

2. Compute the activity rates for the activity cost pools. (Round your answers to 2 decimal places.)

3. The company recently completed a 4,200-square-metre snow removal job at Hometown Hardware—a 82-kilometre round-trip journey from Kenosha’s offices. Compute the cost of this job using the ABC system. (Round your answers to 2 decimal places.)

4. The revenue from the Hometown Hardware job was $485.10 (4,200 square metres at $11.55 per hundred square metres). Prepare a report showing the margin from this job. (Round your answers to 2 decimal places.)
Answer #1
1
Snow removal Travel to Jobs Job support Other Total
Wages 136800 17100 0 17100 171000
Supplies 43500 0 0 0 43500
Snow removal equipment depreciation 19975 0 0 3525 23500
Vehicle expenses 0 28275 0 15225 43500
Office expense 0 0 30150 36850 67000
President’s compensation 0 0 32900 61100 94000
Total cost 200275 45375 63050 133800 442500
2
Activity Cost Pool Activity Rate   
Snow removal 5.00 per hundred square metre =200275/40055
Travel to jobs 2.20 per kilometre =45375/20625
Job support 252.20 per job =63050/250
3
Activity Cost Pool ABC cost
Snow removal 210.00 =42*5
Travel to jobs 180.40 =82*2.20
Job support 252.20
Total cost 642.60
4
Customer Margin—Activity-Based Costing
Sales 485.10
Costs:
Snow removal 210.00
Travel to jobs 180.40
Job support 252.20 642.60
Customer margin -157.50